# IR Bounds on Theories with Spontaneously-Broken Lorentz Symmetry

### v2

@@ Abstract
In nature, some UV features of the dynamics are reflected in IR quantities. In fully relativistic theories, this connection can be probed through the analyticity properties of scattering amplitudes, allowing to understand which IR theories respect the UV assumptions of quantum field theory. The ensuing analyticity bounds can be usually rephrased as the absence of faster-than-light propagation for low-energy excitations. While it is interesting to understand these relations and their IR characterization for theories that have less idealized properties, it is also more difficult to derive analyticity bounds in these cases. For theories that spontaneously break Lorentz symmetry, recent progress was made by considering correlators of conserved currents and their analyticity properties. In this work, we focus on such theories and close the gap from the IR side, deriving bounds from the speed of propagation that are equivalent to the known analyticity bounds. Our bounds require that gapped excitations have a slower speed than gapless ones, at least for momenta that are low with respect to the mass gap. These results suggest a way to study the UV/IR connection in more complex theories.
1 Introduction
@@
Separation of scales in nature allows to describe low-energy (IR) dynamics without knowledge of high-energy (UV) degrees of freedom. At the same time, low-energy interactions encode some of the properties that characterize the high-energy dynamics.
@@
This connection between UV and IR becomes remarkably transparent for theories that in the UV are compatible with unitarity, Lorentz invariance, locality and microcausality – the property that operators commute when space-like separated Eden1971 . These assumptions about the UV theory entail polynomial boundedness and canonical analyticity properties of S-matrix elements as functions of the Mandelstam invariants Bros1964 ; Martin:1969 ; Streater:1989 , leading to bounds on the coefficients of low-energy interactions, see e.g. Pham:1985 ; Adams:2006sv ; deRham:2017avq ; Caron-Huot:2020cmc ; Bellazzini:2020cot ; Tolley2021 ; Sinha:2020win ; Bellazzini2022 ; Arkani-Hamed2021a ; Caron-Huot2022 .
@@
In light of this UV/IR connection, it is interesting to explore whether these boundaries in the space of low-energy Wilson coefficients can be characterized in a synthetic way in terms of IR dynamical features. Such an IR characterization has been explored e.g. in Adams:2006sv ; Camanho:2014apa ; deRham:2021bll ; Serra2022 ; CarrilloGonzalez2022 ; CarrilloGonzalez2023 ; Serra:2023nrn , finding that certain effective field theories (EFTs) that do not respect the analyticity bounds display faster-than-light propagation for excitations around non-trivial backgrounds. These results exposed that the phase velocity captures, at least in some cases, the UV properties that lead to analyticity bounds. In particular, in various examples of Lorentz invariant theories, one can argue (and sometimes prove) that analyticity bounds forbid classical time advances at least on a wide class of backgrounds; see e.g. Arkani-Hamed2021a ; Caron-Huot2022 ; Bellazzini:2022wzv .
@@
This may be phrased in terms of the relation between microcausality and phase velocity. Indeed, the limiting value of the phase velocity at high frequency, a.k.a. the front velocity, should be insensitive to any fixed wavelength background and should therefore fall into the light-cone as dictated by microcausality Brillouin:1960tos ; Diener:1996mj . Therefore, for cases in which the phase velocity does not decrease in the UV, its value at low frequency should always be lower than the speed determined by microcausality [^1] See e.g., Diener:1996mj ; Greene:2022uyf for interesting exceptions in this regard..
@@
Given these results, it is valuable to understand how the same connection between UV and IR becomes manifest in more general contexts, where Lorentz invariance, locality, or other assumptions may be to some extent relaxed, in either or both the UV and the IR. For instance, EFTs describing physical systems like condensed matter or cosmology may carry information about the underlying UV theory, even if the low-energy interactions nonlinearly realize Lorentz. As a matter of fact, analyticity bounds are rooted in the study of propagation of light in a material, see e.g. Keldysh1989TheDF ; Creminelli:2024lhd and have recently been fruitfully extended to more generic theories that spontaneously break Lorentz invariance Creminelli:2022onn . Similarly, less conservative assumptions about the UV may still entail rich analyticity structures and map out boundaries in the space of Wilson coefficients that might be measured in the future, see e.g. Haring:2022cyf ; Buoninfante:2023dyd ; Buoninfante:2024ibt for recent efforts in this direction. In this regard, it would be very interesting to understand whether analyticity bounds and the corresponding UV assumptions have a universal translation in terms of IR dynamical features.
@@
In this work, as a first step towards understanding the IR characterization of analyticity bounds beyond the most conservative assumptions, we focus on the case of an EFT that spontaneously breaks Lorentz symmetry. In this context the analyticity of correlators of conserved currents can be exploited to obtain bounds on the coefficients of low-energy interactions Creminelli:2022onn . We parallel these results by studying how the speed of propagation of low-energy excitations in the EFT is affected by interactions, deriving bounds on the Wilson coefficients in terms of the excitations’ velocity. In this way, we provide the IR analog of the analyticity bounds derived in Creminelli:2022onn , confirming that the velocity of low-energy excitations can meaningfully capture the information contained in the analyticity bounds. As we discuss below, we find that a more general notion of velocity, going beyond the phase velocity, is needed to characterize the IR system in a synthetic way.
@@
In the specific, we consider the EFT of a conformal superfluid coupled to a gauge field in $D=3$ space-time dimensions. These systems are characterized by a non-linear realization of time translations due to a finite charge density phase of a scalar field charged under $U(1)$. This phase breaks the conformal symmetry to spatial rotations and translations plus the non-linearly realized time-translation. The IR dynamics of these systems has been studied extensively in the context of large charge expansion, see e.g. Hellerman:2015nra ; Monin:2016jmo ; Cuomo:2020rgt ; Cuomo:2021qws .
@@
In these systems, the existence and analyticity properties of the S-matrix become nontrivial Hui:2023pxc ; Creminelli:2023kze , requiring one to consider alternative approaches in order to derive analyticity bounds. In Creminelli:2022onn this was done by studying correlators of the conserved $U(1)$ currents, due to their convenient UV behavior and analytic properties. In practice, these correlators were computed by adding a background, non-dynamical $U(1)$ gauge field coupled to the superfluid. While the $U(1)$ symmetry can be easily gauged, it is not simple to study the dynamics of the gauge field and to prospect which UV completions would give rise to the gauged EFT in the broken phase.
@@
In our work, we follow an approach similar in spirit to what was done in Creminelli:2022onn and introduce a $U(1)$ gauge field coupled to the conformal superfluid, with vanishing background field-strength. Differently from Creminelli:2022onn , we consider the gauge field in this IR system as dynamical. While it might be subtle to interpret this IR system as a physical system that is UV completed by a conformal field theory, we find it useful to consider this setup simply as an auxiliary system to extract key information about the conformal superfluid, similarly to how the background gauge field is used in Creminelli:2022onn . In this IR system, we imagine that a background charge density makes the configuration with vanishing background field strength stable. Our point of view follows the natural expectation that a background field may be regarded as an approximation of a dynamical field whose excitations are being neglected. As a matter of fact, the quadratic part of the effective action for the $U(1)$-gauged conformal superfluid allows to consistently describe the excitations of both the scalar and vector fields, even when one writes this action imagining a non-dynamical, background vector field. In this action, a kinetic mixing converts scalar modes into vector modes and viceversa. As we discuss, this kinetic mixing allows to recover the information contained in correlators of conserved currents by examining the propagation of the low-energy modes.
@@
A universal feature of the EFTs for the broken phase is that gapless excitations at zero momentum have soundspeed $c_{s}^{2}=1/2$ in $D=3$. This means that in any UV complete model that has Lorentz symmetry in the UV, the phase velocity of a gapless mode will change from $1/2$ to $1$ as one explores excitations with higher and higher momenta (eventually outside of the regime of validity of the EFT for the broken phase). If the low-energy theory has a mass gap, the phase velocity cannot be used to characterize the system at momenta below the mass gap, as it is divergent at zero momentum. A better notion of velocity is the one associated with the trajectories of rays Boillat:1969slt ; Boillat:1970gw ; Landau ; Sawicki:2024ryt , which coincides with the phase velocity for gapless excitations and has a meaningful non-relativistic limit (vanishing at zero momentum, for gapped excitations). With this, one can formulate a meaningful bound for the EFT by asking that the ray velocity of gapped excitations does not grow too much for momenta below the mass gap. We apply this reasoning to the gauged conformal superfluid, since the kinetic mixing between scalar and vector excitations leads to a mass gap for both scalar and vector modes. In practice we ask the ray velocity to be at most $1/2$ for momenta lower than the mass gap. In this way, the excursion between IR and UV values of ray velocity is entirely captured by the (effectively) gapless, high momentum regime of the theory, as one might expect in analogy with the gapless cases. This IR requirement, as we show, leads to the same bounds found in Creminelli:2022onn studying correlators of conserved currents.
@@
This result makes it clear that the IR meaning of the bounds is not the absence of paradoxical causality violations, as previously thought Adams:2006sv . Indeed, propagation outside the speed $1/2$ soundcone is not in conflict with causality, while it is forbidden by the analyticity bounds. This is consistent with the finding that no causal paradoxes associated with superluminality in the IR can be predicted in the EFT, as discussed in Babichev:2007dw ; Burrage:2011cr ; Kaplan:2024qtf . The work is organized as follows. In Section 2 we describe the model of gauged conformal superfluid and its propagating modes at low frequencies. In Section 3 we present a precise derivation of ray velocity for gapless and gapped modes using the acoustic metric description. In Section 4, we obtain the IR bounds on the EFT of conformal superfluids using the ray velocity, and compare our result with the analyticity bounds of Creminelli:2022onn . We close in Section 5 with a brief discussion. Throughout this paper we use the (,+,,+)(-,+$\dots$,+)( - , + , … , + ) signature convention.
2 Conformal superfluid coupled to a gauge field
@@
We start by considering an effective field theory for a conformal superfluid in $D=3$ coupled to a vector field, as discussed in Ref. Creminelli:2022onn ,
@@
As we outline in Sec. 4.3, such an EFT arises e.g. when integrating out the radial mode of a $U(1)$ charged scalar in a conformal field theory. If the vector field was a dynamical degree of freedom in the UV, then one would expect a purely kinetic term for the gauge field to appear in Eq. (1), for instance derived using the symmetry arguments in Cuomo:2021qws . To keep clear the analogy with the work in Creminelli:2022onn , we will neglect this term and use the above EFT only as an auxiliary IR system to extract information about the conformal superfluid. As we show below, the interactions between scalar and vector already endow the vector with a healthy kinetic term in the broken phase. This theory only makes sense on top of a time-dependent background for the scalar field
@@
which is responsible for the spontaneous breaking of Lorentz invariance below the scale $\mu$, which we assume $\mu>0$ without loss of generality, and which acts as the cutoff of the EFT for the broken phase. This theory enjoys a gauge symmetry
@@
with $i=1,2$ running over the spatial directions. This Lagrangian provides both kinetic terms for $\pi$ and for $A_{\mu}$, which are free of ghost, gradient and tachyonic instabilities for $c_{1}>0$, $b>0$ and $b+d>0$. Therefore, this quadratic Lagrangian and the related Hamiltonian, consistently describe the vector and scalar fields as healthy, dynamical variables. Importantly, there is also kinetic mixing between $\pi$ and $A_{\mu}$ which implies that if one excites either of the two fields, inevitably the other will get excited as well. Due to this, we should study the eigenmodes of propagation.
As discussed in the Introduction, we do not insist that the gauge field should be a physical degree of freedom of the UV conformal theory, and content ourselves with regarding it as part of an IR auxiliary system we use to study the conformal superfluid. Auxiliary or not, we imagine that the IR system describing the broken phase with vanishing electric potential is stable thanks to a suitable homogeneous background charge. This background charge makes possible for the scalar field to get a non-zero charge density while satisfying the Gauss law with zero background electric field. This phase will be stable if one chooses appropriate self interactions for the scalar.
2.1 Scalar gapless mode
It is instructive to first consider the decoupling limit when the vector field is in fact nondynamical (formally $b+d\to\infty$). In this case, the dispersion relation for the canonical scalar $\pi_{c}=\sqrt{\mu c_{1}}\pi$ is obtained as

### v3

# IR side of bounds on Theories with Spontaneously Broken Lorentz Symmetry

@@ Abstract
In nature, some UV features of dynamics are reflected in IR quantities. In fully relativistic theories, this connection can be probed through the analyticity properties of scattering amplitudes, allowing one to understand which IR theories respect the UV assumptions of quantum field theory. The ensuing analyticity bounds can usually be rephrased as the absence of faster-than-light propagation for low-energy excitations. While it is interesting to understand these relations and their IR characterization for theories that have less idealized properties, it is also more difficult to derive analyticity bounds in these cases. For theories that spontaneously break Lorentz symmetry, recent progress was made by considering correlators of conserved currents and their analyticity properties. In this work, we focus on such theories and work to close the gap from the IR side, finding a natural way to express the known analyticity bounds purely in terms of low-energy kinematical quantities. Our analysis shows that the bounds require gapped excitations to have a slower speed than the gapless ones, at least for momenta that are low with respect to the mass gap. These results suggest a way to interpret the UV/IR connection in more complex theories.
I Introduction
@@
Separation of scales in nature allows one to describe low-energy (IR) dynamics without knowledge of high-energy (UV) degrees of freedom. At the same time, low-energy interactions encode some of the properties that characterize the high-energy dynamics.
@@
This connection between UV and IR becomes remarkably transparent for theories that in the UV are compatible with unitarity, Lorentz invariance, locality and microcausality – the property that operators commute when space-like separated [1]. These assumptions about the UV theory entail polynomial boundedness and canonical analyticity properties of S-matrix elements as functions of the Mandelstam invariants [2, 3, 4], leading to bounds on the coefficients of low-energy interactions, see e.g. [5, 6, 7, 8, 9, 10, 11, 12, 13, 14].
@@
In light of this UV/IR connection, it is interesting to explore whether these boundaries in the space of low-energy Wilson coefficients can be characterized in a synthetic way in terms of IR dynamical features. Such an IR characterization has been explored e.g. in [6, 15, 16, 17, 18, 19, 20], finding that certain effective field theories (EFTs) that do not respect the analyticity bounds display faster-than-light propagation for excitations around non-trivial backgrounds. These results exposed that the speed of propagation of low-energy excitations captures, at least in some cases, the UV properties that lead to analyticity bounds. In particular, in various examples of Lorentz invariant theories, one can argue (and sometimes prove) that analyticity bounds forbid classical time advances at least on a wide class of backgrounds; see e.g. [13, 14, 21].
@@
This may be phrased in terms of the relation between microcausality and the kinematics of low-energy excitations. Indeed, the limiting value of the phase velocity at high frequency, also known as the front/signal velocity, should be insensitive to any fixed wavelength background and should therefore fall into the light-cone as dictated by microcausality [22, 23]. Similarly, in the regime where signal propagation can effectively be described as classical trajectories/rays, the group velocity encodes the support of Green’s functions [23]. Therefore, for cases in which group and/or phase velocity do not decrease in the UV, their values at low frequency should be lower than the speed determined by microcausality [^1] See e.g., [23, 24] for interesting exceptions in this regard..
@@
Given these results, it is valuable to understand how the same connection between UV and IR becomes manifest in more general contexts, where Lorentz invariance, locality, or other assumptions may be to some extent relaxed, in either or both the UV and the IR. For instance, EFTs describing physical systems like condensed matter or cosmology may carry information about the underlying UV theory, even if the low-energy interactions nonlinearly realize Lorentz invariance. As a matter of fact, analyticity bounds are rooted in the study of propagation of light in a material, see e.g. [25, 26] and have recently been fruitfully extended to more generic theories that spontaneously break Lorentz invariance [27]. Similarly, less conservative assumptions about the UV may still entail rich analyticity structures and map out boundaries in the space of Wilson coefficients that might be measured in the future, see e.g. [28, 29, 30] for recent efforts in this direction. In this regard, it would be very interesting to understand whether analyticity bounds and the corresponding UV assumptions have a universal translation in terms of IR dynamical features.
@@
In this work, as a first step toward understanding the IR characterization of analyticity bounds beyond the most conservative assumptions, we focus on the case of an EFT that spontaneously breaks Lorentz symmetry. In this context, the analyticity of correlators of conserved currents has been exploited to obtain bounds on the coefficients of low-energy interactions [27]. We examine these results by studying how the propagation of low-energy excitations in the EFT is affected by interactions, expressing the known bounds on the Wilson coefficients in terms of the excitations’ kinematics. In this way, we provide an IR interpretation of the analyticity bounds derived in [27], showing that a specific notion of velocity for low-energy excitations sharply captures the information contained in the analyticity bounds. In fact, our result is that the acoustic vector, defined in terms of the effective geometry that describes propagation at low energy and the momentum (wave-number) co-vector, must lay inside a fixed, momentum-independent, sound-cone. This remarkably simple condition, characterizes synthetically and in a momentum-independent way the analyticity bounds in terms of the IR kinematics.
@@
Specifically, we consider the EFT of a conformal superfluid coupled to a gauge field in $D=3$ space-time dimensions. These systems are characterized by a non-linear realization of time translations due to a finite charge density phase of a scalar field charged under $U(1)$. This phase breaks the conformal symmetry to spatial rotations and translations plus the non-linearly realized time-translation. The IR dynamics of these systems has been studied extensively in the context of large charge expansion, see e.g. [31, 32, 33, 34]. We stress that this choice of EFT solely follows the need for a tractable example for which analyticity bounds have been derived, and is to be considered as a proxy for more complicated situations such as theories relevant for modeling cosmology. In the same vein, our studies here are intended as a first step towards a more complete IR characterization of analyticity bounds.
@@
In these systems, the existence and analyticity properties of the S-matrix become nontrivial [35, 36], requiring one to consider alternative approaches in order to derive analyticity bounds. In [27] this was done by studying correlators of the conserved $U(1)$ currents, due to their convenient UV behavior and analytic properties. In practice, these correlators were computed by adding a background, non-dynamical $U(1)$ gauge field coupled to the superfluid. While the $U(1)$ symmetry can be easily gauged, it is not simple to study the dynamics of the gauge field and to prospect which UV completions would give rise to the gauged EFT in the broken phase.
@@
In our work, we follow an approach similar in spirit to what was done in [27] and introduce a $U(1)$ gauge field coupled to the conformal superfluid, with vanishing background field-strength. Similarly to [27], the gauge field is considered as an auxiliary system to extract key information about the conformal superfluid. While [27] considers the gauge field as fixed background, we keep track of its kinematics. In fact, it is natural to regard the background field as an approximation of a dynamical field whose excitations are being neglected. While it might be subtle to interpret this IR system as a physical system that is UV completed by a conformal field theory, we imagine that a background charge density makes the configuration with vanishing background field strength stable. As a matter of fact, the quadratic part of the effective action for the $U(1)$-gauged conformal superfluid allows to consistently describe the excitations of both the scalar and gauge fields, even when one writes this action imagining a non-dynamical, background gauge field. In this action, a kinetic mixing converts scalar modes into modes of the gauge field and viceversa. As we discuss, this kinetic mixing allows one to recover the information contained in correlators of conserved currents by examining the propagation of the low-energy modes.
@@
A universal feature of the EFTs for the broken phase is that gapless excitations at zero momentum have soundspeed $c_{s}^{2}=1/2$ in $D=3$. This means that in any UV complete model that has Lorentz symmetry in the UV, the phase velocity of a gapless mode will change from $1/2$ to $1$ as one explores excitations with higher and higher momenta (eventually outside of the regime of validity of the EFT for the broken phase). If the low-energy theory has a mass gap, the phase velocity cannot be used to characterize the system at momenta below the mass gap, as it is divergent at zero momentum. In these cases, one can use a different notion of velocity associated with wave packets [37, 38, 39, 40], the group velocity. This coincides with the phase velocity for several gapless systems, but keeps being well-defined for gapped excitations at zero momentum. Here we find that neither of these notions of velocity can be used to express the known analyticity bounds in a sharp, momentum-independent way. This difficulty prompts us to instead study directly the propagation through the effective acoustic metric seen by IR excitations. In this language, an acoustic vector can be naturally defined, leading to a notion of velocity which is closely related to both the phase and the group velocity, but that differs from these two in general. (In fact, as we will show, this new velocity matches the phase velocity for gapless excitations, while it coincides with the group velocity in other cases.) From this point of view, we find that the analyticity bounds are encoded in a momentum-independent condition on the acoustic vector of IR excitations and its associated velocity. Physically, this condition requires gapped excitations to have a lower velocity than the gapless ones, for momenta below the mass gap. We apply this reasoning to the gauged conformal superfluid, since the kinetic mixing between scalar and gauge-field excitations leads to a mass gap for all the physical modes. In practice, we ask the velocity associated to the acoustic vector to be at most $1/2$ for momenta lower than the mass gap. In this way, the excursion between IR and UV values of the velocity is entirely captured by the (effectively) gapless, high momentum regime of the theory where the velocity defined by the acoustic vector asymptotes to the phase velocity, as one might expect in analogy with the gapless cases. This IR requirement, as we show, leads to the same bounds found in [27] studying correlators of conserved currents.
@@
This result makes it clear that the IR meaning of the analyticity bounds is not the absence of paradoxical causality violations, as previously thought [6]. Indeed, propagation outside the speed $1/2$ soundcone is not in conflict with causality, while it is forbidden by the analyticity bounds. This is consistent with the finding that no causal paradoxes associated with superluminality in the IR can be predicted in the EFT, as discussed in [41, 42, 43]. The work is organized as follows. In Section 2 we describe the model of gauged conformal superfluid and its propagating modes at low frequencies. In Section 3 we discuss the IR kinematics for gapless and gapped modes using the acoustic metric description, introducing and motivating various natural notions of velocity. In Section 4, we consider the analyticity bounds of [27] on the EFT of conformal superfluids, expressing them as a momentum-independent condition on the acoustic vector of low-energy excitations. We close in Section 5 with a brief discussion. Throughout this paper we use the (,+,,+)(-$\dots,+$) signature convention. We refer to vectors $v=v^{\mu}\partial_{\mu}$ and co-vectors $w=w_{\mu}dx^{\mu}$, respectively, in terms of their contravariant $v^{\mu}$ and covariant $w_{\mu}$ components, with $\{\partial_{\mu}\}$ and $\{dx^{\mu}\}$ being the corresponding coordinate basis of the tangent and co-tangent spaces.
II Conformal superfluid coupled to a gauge field
@@
We start by considering an effective field theory for a conformal superfluid in $D=3$ coupled to a gauge field, as discussed in Ref. [27],
@@
As we outline in Sec. IV.4, such an EFT arises e.g. when integrating out the radial mode of a $U(1)$ charged scalar in a conformal field theory. If the gauge field was a dynamical degree of freedom in the UV, then one would expect a purely kinetic term for the gauge field to appear in Eq. (1), for instance derived using the symmetry arguments in [34]. To keep clear the analogy with the work in [27], we will neglect this term and use the above EFT only as an auxiliary IR system to extract information about the conformal superfluid. As we show below, the interactions between scalar and gauge field already endow the latter with a healthy kinetic term in the broken phase. This theory only makes sense on top of a time-dependent background for the scalar field
@@
which is responsible for the spontaneous breaking of Lorentz invariance below the scale $\mu$, which we assume as $\mu>0$ without loss of generality, and which acts as the cutoff of the EFT for the broken phase. This theory enjoys a gauge symmetry
@@
with $i=1,2$ running over the spatial directions. This Lagrangian provides both kinetic terms for $\pi$ and for $A_{\mu}$, which are free of ghost, gradient and tachyonic instabilities for $c_{1}>0$, $b>0$ and $b+d>0$. Therefore, this quadratic Lagrangian and the related Hamiltonian, consistently describe the gauge and scalar fields as healthy, dynamical variables. Importantly, there is also kinetic mixing between $\pi$ and $A_{\mu}$ which implies that if one excites either of the two fields, inevitably the other will get excited as well. Due to this, we should study the eigenmodes of propagation.
As discussed in the Introduction, we regard the gauge field as part of an IR auxiliary system we use to study the conformal superfluid. We imagine that the IR system describing the broken phase with vanishing electric potential is stable due to a suitable homogeneous background charge. This background charge makes it possible for the scalar field to get a non-zero charge density while satisfying the Gauss law with zero background electric field. This phase will be stable if one chooses appropriate self interactions for the scalar.
The coefficients $c_{2}$ and $c_{3}$ of the Higher-Dimensional (HD) operators have been shown to be restricted by analyiticity to satisfy [27]
$\frac{c_{2}}{b+d}\left(1-\xi^{2}\right)-\frac{c_{3}}{b+d}\geq-\frac{\left(1-\frac{\xi^{2}}{2}\right)^{2}}{\xi^{2}}\,$, (7)
where $\xi\in[0,1]$. We are interested in expressing this bound in terms of IR kinematic quantities, for which we proceed to study the propagating degrees of freedom.
II.1 Scalar gapless mode
It is instructive to first consider the decoupling limit when the gauge field is in fact nondynamical (formally $b+d\to\infty$). In this case, the dispersion relation for the canonical scalar $\pi_{c}=\sqrt{\mu c_{1}}\pi$ is obtained as